An apartment building contains 12 units consisting of one- and two-bedroom apartments that rent for $360 and $450 per month, respectively. When all units are rented, the total monthly rental is $4,950. What is the number of two-bedroom apartments?
A. 3
B. 4
C. 5
D. 6
E. 7

Answer by rfer(16322)

x+y=12 y=12-x
360x+450y=4950
360x+450(12-x)=4950
360x+5400-450x=4950
-90x=-450
x=5
y=12-5=7 two bed